OpenStudy (anonymous):

PLEASE HELP! MEDAL WILL DEFINIETLY BE GIVEN!! Express answer in exact form. Find the area of the larger segment whose chord is 8" long in a circle with an 8" radius. (Hint: A chord divides a circle into two segments. In problem 1, you found the area of the smaller segment.)

OpenStudy (anonymous):

In problem 1 the smaller segment was-->\[A=(32/3 \pi - 16 sq rt 3) inches ^2\]

OpenStudy (ranga):



OpenStudy (ranga):

Oh, then subtract the smaller segment area from the area of the entire circle to get the area of the larger segment.

OpenStudy (anonymous):

What's the area of the entire circle

OpenStudy (ranga):

Area of a circle = pi(r^2) r = 8 Area = pi(8^2) = 64pi

OpenStudy (anonymous):

So the answer would be...

OpenStudy (ranga):

Area of the larger segment = Area of circle - Area of smaller segment = 64pi - [ (32/3)(pi) - 16sqrt(3) ] Just simplify the above.

OpenStudy (anonymous):

Ohhh otay :j

OpenStudy (anonymous):

Thank you soo much!!! If I needed some more help would you help me please?